### v1

# Modeling phase transformations in Mn-rich disordered rocksalt cathodes with charge-informed machine-learning interatomic potentials

Abstract
Mn-rich disordered rocksalt (DRX) cathode materials exhibit a phase transformation from a disordered to a partially disordered spinel-like structure (-phase) during electrochemical cycling. In this computational study, we used charge-informed molecular dynamics with a fine-tuned CHGNet foundation potential to investigate the phase transformation in LixMn0.8Ti0.1O1.9F0.1. Our results indicate that transition metal migration occurs and reorders to form the spinel-like ordering in an FCC anion framework. The transformed structure contains a higher concentration of non-transition metal (0-TM) face-sharing channels, which are known to improve Li transport kinetics. Analysis of the Mn valence distribution suggests that the appearance of tetrahedral Mn2+ is a consequence of spinel-like ordering, rather than the trigger for cation migration as previously believed. Calculated equilibrium intercalation voltage profiles demonstrate that the -phase, unlike the ordered spinel, exhibits solid-solution signatures during the 0-TM to Li conversion reaction. A higher Li capacity is obtained than in the DRX phase. This study provides atomic insights into solid-state phase transformation and its relation to experimental electrochemistry, highlighting the potential of charge-informed machine learning interatomic potentials for understanding complex oxide materials.
I Introduction
The increasing demand for sustainable energy storage has driven the development of high-energy-density and cost-effective cathode materials for rechargeable Li-ion batteries [1, 2, 3]. Disordered rocksalt (DRX) cathodes using manganese (Mn) as the earth-abundant redox-active cation have the potential to scale Li-ion energy storage to several TWh/year [4]. Recent studies demonstrate that initiating a Mn-rich disordered rocksalt structure undergoes an in-situ transformation into a partially disordered spinel-like phase (termed as -phase) upon electrochemical cycling, resulting in increased capacity and rate capability for cathode designs [5, 6, 7, 8, 9]. This approach mitigates the typical low capacity in other DRX cathodes by promoting a new voltage feature around 4 V and an extended 3 V plateau. Moreover, in contrast to conventional LixMn2O4 spinel cathodes, which suffer from the two-phase reactions leading to inhomogeneity-induced stresses and mechanical degradation between , the -phase maintains a single (solid-solution) phase reaction throughout the cycling process. Cai et al. [9] demonstrates that such a series of -phase material exhibits an impressive capacity of over 280 mAh/g and an energy density exceeding 800 Wh/kg while maintaining minimal voltage decay even after extensive cycling.
The spinel is distinct from the disordered rocksalt based on cation ordering in an FCC anion framework. Figure 1a–1c presents the cation orderings in a conceptualized graph representation as a projection of the 3-dimensional FCC structure shown in Fig. 1c, including the disordered rocksalt (Fmm), spinel (Fdm), and partially disordered spinel (-phase) structures. The green and purple spheres represent Li and Mn ions, respectively, occupying the octahedral sites (, denoted as ) of the FCC rocksalt framework. For simplicity, the stoichiometric LiMnO2 is used for illustration in Fig. 1a-c. The graph edges connect the sites to the interstitial tetrahedral sites (, denoted as ), which are represented by white dashed circles in Fig. 1. In the spinel LiMnO2 ordering, the octahedral sites are categorized into and sites as shown in Fig. 1b. All sites are occupied by Li, while all sites are occupied by Mn. The tetrahedral site neighbored by four Li is denoted as an site in a spinel. When the spinel is delithiated to Li0.5MnO2, the sites become vacancies and the sites are all occupied by Li. More generally, the tetrahedral site neighbored by four Li is also termed a 0-TM channel. Urban et al. [10] demonstrated that long-range connected 0-TM percolating channels are essential for facile Li diffusion in an FCC framework, with the spinel ordering having the highest amount of 0-TM channels in theory. The partially disordered spinel is conceptually illustrated in Fig. 1c as an intermediate state between the spinel and disordered ordering.
The phase transformation from DRX to -phase is related to the transition metal (TM) migration of the mobile Mn ion when neighboring cation vacancies are created [11]. The TM can take the pathway connected by the octahedral-tetrahedral-octahedral (--) sites in an FCC framework. For the migration process, it is typically concluded that Mn migration is strongly coupled with its charge state, Mn4+ is generally considered immobile, while Mn3+ and Mn2+ are considered more mobile according to the nudged-elastic band calculations [12, 13, 14]. However, simulating these charge-coupled dynamics has long been challenging, as the required time and length scales are inaccessible to conventional ab initio methods. Previously, cluster expansion (CE) methods have been shown successfully as a surrogate model of density functional theory (DFT) calculations for modeling the configurational thermodynamics in complex oxides via Monte Carlo (MC) simulation at scale [15, 16, 17]. Nonetheless, the limited expressibility and coarse-grained nature of these models render them less adequate for capturing the complex charge-coupled dynamical processes involved in TM migrations [18, 9]. The advent of machine learning interatomic potentials (MLIPs) has opened new avenues for addressing these problems by enabling large-scale and nano-second-long simulations with ab-initio quantum accuracy [19, 20, 21]. By using magnetic moments as a proxy for atomic charge, Deng et al. [21] recently simulated a phase transformation from an orthorhombic to a spinel-like phase in Li0.5MnO2 with charge-informed MD. It identified the timescale discrepancies between charge disproportionation and ion hops, suggesting that the emergence of Mn2+ is correlated to long-range spinel-like ordering but not solely to the local hopping into the tetrahedral occupancy.
In this study, we extend the computational study of the phase transformation in LixMn0.8Ti0.1O1.9F0.1 (LMTOF), which is a Mn-rich DRX cathode. We parametrized a CHGNet-MLIP using r2SCAN-DFT calculations for various atomic configurations to sample the potential energy surface in this chemical space. The fine-tuned CHGNet-MLIP was used to implement charge-informed MD simulations starting from a disordered phase in Li0.6Mn0.8Ti0.1O1.9F0.1. The MD simulations revealed a phase transformation from a disordered rocksalt to a spinel-like () phase. From the magnetic moment predictions, we demonstrate that the Mn migration is mainly contributed by Mn3+ at the beginning, while Mn2+ collectively emerges as the spinel-like ordering forms. Furthermore, we investigated the impact of this partial disorder on the electrochemical intercalation voltage profiles based on the structures obtained from the MD simulations. Our work provides insights into the application of MLIPs for studying complex energy materials and offers a mechanistic understanding of the partial cation-disorder effect on intercalation electrochemistry at the atomic scale.
II Methods
II.1 Machine learning interatomic potential
To efficiently compute the energy and interatomic forces, we employed the Crystal Hamiltonian Graph Neural Network (CHGNet) as a surrogate model for DFT calculations [21]. CHGNet is a graph-neural-network-based MLIP that utilizes message-passing layers to propagate atomic information via a set of nodes connected by edges [22]. After message-passing layers, the node features are linearly projected onto magnetic moments by , and the total energy is calculated as a sum of non-linear node projections by over all atoms with an additional message-passing layer
The nodes are regularized by DFT magnetic moments to carry rich information about both local ionic environments and to serve as a proxy for the atomic charge. The forces and stress are calculated via auto-differentiation of the total energy with respect to the atomic Cartesian coordinates and strain:
This approach allows CHGNet to capture the complex relationships between charge distributions and energetics with given atomic configurations.
II.2 Potential energy surface sampling
While CHGNet [21] as a pre-trained foundation potential achieves reasonable performance in predicting materials stability on out-of-distribution datasets [23], recent benchmarks observed a systematic softening of the ML predicted potential energy surface (PES) compared to the DFT ground truth, particularly for atomic configurations from MD simulations with high energy [24, 25]. To improve simulation accuracy, we performed multi-faceted sampling of the PES for the Li–Mn–Ti–O–F chemical space [26], as follows:
Exisiting dataset – We collected the compounds in the Materials Project database within the Li–Mn–Ti–O–F chemical space that have decomposition energy (energy above the convex hull, ) lower than 0.1 eV/atom as a general coverage of such chemical space [27]. Furthermore, we adopted the structures from the training dataset of previously reported cluster expansion (CE) models for DRXs containing Mn2+/Mn3+ as redox-active species (DRX-CE) [28, 29]. All these structures in the DRX-CE dataset have rocksalt lattices but different cation and anion orderings. These structures were relaxed by DFT calculations. The relaxed structures with rocksalt frameworks were used to fit a cluster expansion Hamiltonian (lattice model fitting, see Appendix B). The relaxation trajectories from structures from the two datasets were used to fine-tune a Trial CHGNet.
Partially disordered structures – Given that existing datasets do not include atomic configurations specifically representing partially disordered structures with spinel-like orderings, we developed a custom sampling procedure to generate these configurations. (1) Monte Carlo (MC) simulations were employed to obtain the disordered rocksalt atomic configurations equilibrate at finite temperatures, which were generated in supercells of the primitive spinel cell (see Appendix B for details).
Given these disordered structures, we compared the occupancy of the spinel supercell and the disordered rocksalt in such a supercell lattice. The octahedral cation sites are categorized into and sites. Under this classification, a random distribution of TM yields the disordered structure, while an increased number of TM occupations on sites indicates more spinel-like ordering.
To generate different degrees of spinel order, we applied site exchanges between the two nearest octahedral sites to transform the disordered structure towards the spinel ordering through either (a) an exchange between a TM and a Li via a Li3M tetrahedron; or (b) an exchange between a TM and a Li. The (a) step is to create more spinel-like ordering and the (b) step is to exchange the cation ordering within the sites to potentially create more available Li3M tetrahedron. We assume that TM migration occurs through the -- path with no other TM neighbors in the delithiated states (i.e., through the Li3M tetrahedron, Fig. 3). By iteratively applying site exchanges within different Li3M tetrahedra, one can obtain a collection of structures ranging from 0% (disordered) to nearly 100% spinel ordering.
TM migration pathways – Since TM migration is a rare, high-energy-barrier event [9], especially for unfavorable valence states like Mn4+, we implemented a passive sampling approach to effectively model the diffusion kinetics with the following steps:
For each site exchange step via the Li3M tetrahedron, three different states are generated as A (initial state, TM occupied on site), B (intermediate state, TM occupied on site), and C (final state, TM occupied on site), which are shown in Fig. 3. The other neighboring Li atoms are removed to create vacancies, for which we assume no face-sharing between the migrating TM and cations. The structures in the A/B/C states were relaxed using the Trial-CHGNet fitted with structures from the existing datasets. After structure relaxation, five images interpolated along A-B (- path) and B-C (- path) are evaluated with static DFT calculations, analogous to the intermediate states for nudged elastic band (NEB) calculations. Furthermore, we follow the same approach to generate the direct migration from the - path (A-C) for labeling high-energy and unfavorable states.
In addition, we considered two different delithiation levels for sampling TM migration pathways: (a) A dilute level, where only three Li ions are removed to create a diffusion channel for TM migration (e.g., a pristine supercell structure with Li44Mn32Ti4O76F4 yields a delithiated structure with Li41Mn32Ti4O76F4). This level allows for the investigation of local TM hopping energetics. (b) A medium delithiation level, where 0.5 Li per formula unit (f.u.) is removed from the structure to create delithiated states (e.g., Li ions are randomly removed to create a structure with Li24Mn32Ti4O76F4). This level allows the sampling of delithiated states, representing the charging process in a battery. By sampling partially disordered structures at these two delithiation levels, we include a more comprehensive description of the cation ordering energetics, TM hop, and relevant local environments in which TM hopping occurs at the composition that is energetically favorable for spinel formation (i.e., the composition close to Li0.5TMO2).
By adding the static DFT calculations of these sampled atomic configurations to the relaxation trajectories, another MLIP was parameterized (denoted as Sample-CHGNet) to perform MD simulations for enhanced sampling.
MD trajectories – To further sample the thermally excited atomic configurations at finite temperatures, we employed Sample-CHGNet to perform molecular dynamics (MD) simulations with disordered structures across various compositions. These structures were first generated in the pristine (lithiated) state from the CE-MC simulations at K using supercells of and formula units (f.u.), including Li1.2Mn0.5Ti0.3O1.9F0.1, Li1.2Mn0.6Ti0.2O1.8F0.2, Li1.1Mn0.8Ti0.1O1.9F0.1, and Li1.1Mn0.7Ti0.2O2.0. For each disordered structure, three different delithiated levels (removing 0.2/0.4/0.6 Li per f.u. from the pristine structure) were generated to initiate the MD simulations using the Sample-CHGNet. The MD simulations were run at K for ns with a time step of fs under NVT ensembles. From each MD trajectory, 200 structures were randomly sampled, and their energies were calculated using static DFT and added to the training set. In addition to the training set, we further ran CHGNet-MD simulations in NVT ensembles at K using Sample-CHGNet with a supercell of Li44-xMn32Ti4O76F4 ( f.u.). These trajectories were further evaluated by r2SCAN-DFT static calculations and were set as the test set.
The final CHGNet for production (denoted as Product-CHGNet) was fine-tuned using all structures in the training set (including 88,852 structures) starting from the pretrained model. It achieved a convergence of 2 meV/atom in energy, 68 meV/Å in force, 0.086 GPa in stress, and 0.019 in magnetic moment for the test error. A summary flowchart is illustrated in Fig. 2 to describe the whole procedure.
III Results
To investigate the structural evolution, we performed charge-informed molecular dynamics from the disordered structure to monitor the phase change to the -phase. The initial (disordered) structure was generated by sampling from the equilibrium ensemble of CE-MC simulations at K. The MC structure represents the pristine (lithiated) composition (Li1.1Mn0.6Ti0.1O1.9F0.1) with a supercell containing 320 atoms. Based on the pristine structure, a delithiated structure (LiMn0.6Ti0.1O1.9F0.1, where represents cation vacancy) was then obtained by randomly removing Li ions and relaxing the delithiated configuration using the Product-CHGNet with a force convergence criterion of eV/Å. After relaxation, MD simulations were subsequently performed for ns with a time step of fs under the NVT ensemble at K. The MD trajectories were collected for subsequent structural analysis.
III.1 Structure ordering analysis
Figure 4a presents the simulated X-ray diffraction (XRD) patterns derived from the structures within the time interval of to ns. To analyze the evolution of cation ordering on the FCC lattice, the atomic positions from the MD snapshots were refined to their nearest ideal Wyckoff sites. The gray dashed lines in Fig. 4a denote the characteristic peaks of spinel ordering (18∘, 35∘, etc.). At the onset of the MD simulation ( ns), there are no obvious peaks indicating spinel ordering. After approximately ns, the intensity of the spinel peaks increases and becomes more significant in subsequent time scales as the MD simulation proceeds, suggesting the formation of the spinel-like -phase.
Figure 4b illustrates the energy landscape across the simulated MD trajectories. The black points denote the potential energies, while the colored points represent the total energies. The energies are depicted relative to their values at the initial state. A notable decrease in energy (approximately 60 meV/anion) is observed for the -phase following a ns MD simulation. This energy reduction indicates that spinel-like cation ordering is thermodynamically favored at the composition Li0.6Mn0.8Ti0.1O1.9F0.1 when sufficient cation vacancies are present. This thermodynamic preference for spinel-like ordering aligns with previous computational studies that demonstrated the stability of spinel structures in Li-Mn-O systems, such as Li0.5MnO2 [30].
To study the change in local cation ordering, we analyzed the short-range order (SRO) indicated by the tetrahedron occupancy as a function of MD simulation time. Figure 4c presents the occurrence of 0-TM (solid green line) and 1-TM tetrahedral (solid purple line) channels in the MD structure, as a ratio to all interstitial tetrahedral sites. The 0-TM channel has a low migration barrier that kinetically favors fast Li-ion transport [31]. To compare with disordered phases, the dashed green lines in Fig. 4c indicate the random limit of the 0-TM channel for two Li contents ( for Li1.0 and for Li1.1 at the fully lithiated level) for reference. At the beginning of the MD simulation ( ns), the occurrence of 0-TM channels is approximately 0.03. This value is significantly lower than that of the random limit of Li1.1TMO2 (0.09), where the value is estimated based on a rocksalt structure with 0.1 excess Li and fully random cation ordering. The low value of 0-TM content in the initial structure indicates reduced percolating Li content and suggests that SRO is unfavorable for long-range Li percolation, which is consistent with previous predictions in Mn-Ti-based DRXs [28, 29]. As the phase transformation progresses, the occurrence of 0-TM channels increases, reaching approximately 0.06 after ns of MD simulation. This rise in the number of 0-TM channels agrees with the high 0-TM content in the spinel structure, indicating the emergence of spinel-like ordering.
III.2 Charge distribution with ion rearrangement
To understand the correlation between Mn migration and its heterovalent state [9, 32], we inferred atomic charge states from the magnetic moments predicted by CHGNet. Figure 5a presents the site-wise distribution of magnetic moments for Mn ions. The heterovalent states of Mn are categorized as follows: Mn, Mn, and Mn. At the beginning of the MD simulation, Mn predominantly exists in valence states of +3 and +4. As the phase transformation progresses, the concentration of Mn2+ gradually increases, as shown by the increased frequency of the highlighted circles in Fig. 5a, particularly after the emergence of spinel-like ordering from 0.6 to 2.0 ns.
For a detailed description of Mn migration conditioned on its valence state, the occupancy of Mn and Mn per f.u. is plotted in Fig. 5b as an indication of hopping events to tetrahedral sites. Although previous nudged elastic band calculations reveal that Mn2+ has a lower migration barrier [11, 9], the atomic charges predicted by CHGNet show that most of the tetrahedral Mn are Mn from to ns during the MD simulation. This suggests that Mn3+ may migrate without the need for charge disproportionation () as proposed by Reed et al. [11]. The rise in Mn appears after around ns, which corresponds to the formation of spinel characteristic peaks (18∘, 35∘) in Fig. 4a. This evidence suggests that Mn3+ migration initiates the transformation before long-range spinel-like order is established. The subsequent emergence of tetrahedral Mn2+ appears to be a consequence of this long-range ordering (ns-scale) rather than a prerequisite for local cation hops (ps-scale). The simulation results are consistent with previous MD simulations of the LixMnO2 phase transformation revealed by Deng et al. [21] using the GGA-based MLIP.
III.3 Intercalation voltage profile of -phase
To evaluate the effect of partial disorder on the electrochemical performance of the transformed structure from MD, we computed the intercalation voltage profiles using Product-CHGNet. We focused on two characteristic reactions in an ordered spinel LiMnO: (1) a 0-TM conversion reaction at a lower voltage, where the occupancies of Li ions are transformed from the octahedral sites to the tetrahedral sites collectively:
and (2) a conversion at a higher voltage, where Li is intercalated on the site. It is important to note that Eq. (3) is defined for an independent 0-TM channel (i.e., one that does not share Li with another tetrahedral channel). In an ordered LiMnO2 spinel, each Lioct is face-sharing with two 0-TM channels, where removing one Li from such a 0-TM channel results in a high defect energy [18]. The delithiation is, therefore, a collective conversion that exhibits a two-phase reaction between Li0.5MnO2 and LiMnO2 with a characteristic 3 V voltage plateau [33].
Using Eq. (3), we enumerated different Li-vacancy configurations with fixed TMs and an anion framework (topotactically) at each (de)lithiation level based on the MD structure. The Product-CHGNet was used to perform structure relaxation for both lithiation and delithiation paths (see Appendix C for details). The energies of these (de)lithiated structures were used to construct the convex hull and the corresponding voltage profile as shown in Fig. 6a and Fig. 6b (see Methods).
For the region with high Li concentration (), the 0-TM to conversion exhibits nearly solid-solution behavior, as evidenced by the convex hull being connected by several ground states in Fig. 6a. Particularly, for the early stage of delithiation (), the voltage exhibits a flat plateau with a minor slope, indicating that the delithiation through Eq. (3) occurs in similar local chemical environments and exhibits similar chemical potentials (voltages). We found that the voltage at is around 2.40 V, which is slightly lower than the experimental value. This discrepancy is likely attributable to the residual self-interaction error common in DFT functionals, even for r2SCAN, when a Hubbard U correction is not applied [34, 35]. Although the SCAN-based functional removes more self-interaction than previous LDA and GGA functionals [36], it does not fully remove it and underestimates the intercalation voltage, which is consistent with previous reports [18, 17].
Upon reaching , lithium occupies all tetrahedral 0-TM channels (Li) in addition to the octahedral Li not in the 0-TM channels. Two competing delithiation processes are available, involving either the removal of Li or Li. We computed both pathways and found that the low-energy configurations near the convex hull within the range correspond to the removal of Li, while the high-energy configurations (represented by dark red markers in Fig. 6a) correspond to the removal of Li. These results indicate that even in a structure with partial spinel ordering, Li has a preference for delithiation before Li. In contrast, the delithiation in spinel LixMnO2 for entirely corresponds to the removal/insertion of Li with a V plateau. We also observe a voltage increase of V around (the end of 0-TM to Li conversion). The magnitude is not as large as that in spinel LiMnO2 ( V) or in the spinel-ordered LixMn0.8Ti0.1O1.9F0.1 (see Fig. 7c), suggesting the absence of two-phase reactions in the -phase and less inhomogeneity that could lead to battery performance degradation.
To further demonstrate the effect of partial disorder, we evaluated the intercalation voltage profiles of Li1.1Mn0.8Ti0.1O1.9F0.1 in both the DRX and spinel phases for comparison. The DRX structure was selected from the structure derived from the CE-MC simulations (i.e., the configuration at s in the MD simulation). The spinel Li1.1Mn0.8Ti0.1O1.9F0.1 was generated through another CE-MC simulation without exchanges between the and sites, where all sites are occupied by Li1.0 and sites are occupied by Mn0.8, Ti0.1, and Li0.1. The MC simulation allows site exchange within sites and anion sites, which is designed to capture the underlying SRO of anions and cations in DRXs (e.g., Li-F SRO) [37]. The intercalation at high Li content in both DRX and spinel phases only included the 0-TM conversion reaction as described in Eq. (3).
Figure 7a and 7b illustrate the formation energy convex hull and the voltage profile of the DRX phase, with the -phase plotted as a red line for comparison. Due to the insufficient number of 0-TM channels in the DRX structure, a relatively limited Li capacity from the 0-TM conversion is achieved (approximately 0.3 per f.u., as denoted by the green dashed line in Fig. 7a and 7b. The absence of spinel-like ordering results in a profile marked by a mildly steep increase as delithiation progresses. Notably, the overall voltage profile between of the DRX phase is higher than that of the -phase, suggesting a restricted amount of accessible capacity within an operational voltage window, particularly the capacity around V [9, 8]. In contrast, Figure 7b depicts the calculated results for spinel LixMn0.8Ti0.1O1.9F0.1. The voltage profile clearly displays a two-phase reaction between Li1.1 and Li0.6 with a linear convex hull in Fig. 7c. Due to the absence of disorder, the spinel phase exhibits a lower average voltage with a two-phase reaction between , with a voltage step more than 1 V presented at the end of 0-TM conversion. For , the intercalation corresponds to Li, which shows a similar high-voltage feature as the -phase. The effect of cation disorder on the spinel structure demonstrated in our simulations is in full agreement with the previous findings on the cation disorder effects in LiMnO2 spinel by Chen et al. [18].
IV Discussion
In the evolving landscape of modern battery materials, the complexity of components and configurations exemplified by DRX cathode materials necessitates advanced computational models. Particularly, Mn-rich DRXs exhibit a phase transformation to a partially disordered spinel phase (-phase) via in situ electrochemical cycling [9, 8, 38] or chemical delithiation [39]. This transformation is challenging for conventional modeling approaches to capture, particularly the kinetic pathways of TM migration in such a complex oxide. While ab initio molecular dynamics could theoretically address this issue, its computational cost renders it impractical for probing large systems or long timescales. Alternatively, kinetic Monte Carlo (kMC) methods with CE models have been proposed to model cation ordering changes in both cathodes and solid electrolytes [40, 41]. However, parameterizing CEs for disordered oxides/oxyfluorides and accurately calculating migration barriers in diverse local environments remains a non-trivial endeavor [42].
To address these challenges, we fine-tuned a charge-informed foundation MLIP with sampled r2SCAN-DFT calculated energies, forces, stresses, and magnetic moments. The multi-step PES sampling captures the complex interactions from diverse chemical environments and heterovalent states of TMs. Our MD simulations using the fine-tuned CHGNet successfully captured the phase transition from a disordered to a spinel-like configuration in Li0.6Mn0.8Ti0.1O1.9F0.1.
The simulations reveal that the formation of the -phase promotes the occurrence of 0-TM channels, which play a crucial role in facilitating long-range Li-ion transport, whereas the DRX phase exhibits limited availability of these channels. This result agrees with the improved high-rate performance observed in experiments [9, 8, 39]. By analyzing the electrochemical intercalation voltage profiles of the -phase and comparing them to those of the DRX and spinel phases, two conclusions can be drawn from the simulation results: (1) The presence of partial cation disorder in the -phase effectively eliminates the two-phase reaction, as evidenced by the absence of a sharp voltage increase after Li0.5 removal (see Fig. 7c). (2) The introduction of spinel-like ordering increases the 0-TM occurrence (as indicated by the increased number of 0-TM channels in Fig. 4c) and thus expands the available Li capacity from the 0-TM to Li conversion reaction around 3 V (denoted by the green dashed line in Fig. 6b). The Li capacity from 0-TM to Li conversion is even higher than that of the spinel in Fig. 5c since the 0-TM channels in partially disordered phases are less correlated with each other compared to the spinel, where each Li is shared by two 0-TM channels.
Although this study represents an integrated investigation of phase transformation in complex oxides combining MC simulations for configurational thermodynamics and MLIP-MD simulations for kinetic modeling, several limitations were identified that could be further addressed in future studies. For example, while our MLIP enables nanosecond-scale or million-atoms simulations with ab initio accuracy [43], many important physical processes—such as the formation of anti-phase boundaries in the -phase [39] or phase segregation in LixFePO4 cathodes [44]—occur on timescales still beyond this reach. Achieving ab initio accuracy while enabling simulations beyond typical MD timescales (such as using generative models for dynamic trajectories) holds promise for investigating these processes [45, 46, 47]. Regarding model fine-tuning, recent advances such as active learning with Bayesian force fields [48], stratified sampling from MLIP-driven MD [49], and diffusion-based generative models for mapping reaction pathways [50] provide more efficient sampling techniques beyond the passive sampling implemented in this work, which could enhance MLIP training accuracy by exploring the PES more effectively.
In summary, we demonstrated the application of charge-informed MLIP to model the complex phase transformation from DRX to spinel-like -phase in Li-Mn-Ti-O-F systems. The MLIP-MD simulations revealed the mechanism of spinel-like ordering formation and its effect on electrochemical properties. As computational capabilities and new methodologies emerge–particularly in accelerating MLIP simulations and in generative modeling–these advances will pave the way for more accurate and efficient modeling of phase transformations in next-generation energy storage materials.
V Acknowledgments
This work was primarily supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC0205CH11231 (Materials Project program KC23MP), and the Assistant Secretary for Energy Efficiency and Renewable Energy, Vehicle Technologies Office, under the Advanced Battery Materials Research (BMR) Program of the US Department of Energy (DOE) under contract No. DE-AC0205CH11231. The computational modeling in this work was supported by the computational resources provided by the Extreme Science and Engineering Discovery Environment (XSEDE), supported by National Science Foundation grant number ACI1053575; the National Energy Research Scientific Computing Center (NERSC); and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory. The authors thank Zijian Cai for the valuable discussions.
Appendix A DFT Calculations
All DFT calculations were performed with the VASP package using the projector-augmented wave method [51, 52], a plane-wave basis set with an energy cutoff of 680 eV, and a reciprocal space discretization of 25 k-points per Å-1. The calculations were converged to eV in total energy for electronic loops and 0.02 eV/Å in interatomic forces for ionic loops. The regularized strongly constrained and appropriately normed meta-GGA exchange-correlation functional (r2SCAN) [36, 53] was used with consistent computational settings as MPScanRelaxSet [54].
Appendix B Cluster Expansion & Monte Carlo
The cluster expansion (CE) method [55] is used to study the configurational thermodynamics of materials in which sites can be occupied by multiple cations, and has been applied to study the Li-vacancy intercalation chemistry in layered and disordered materials [56, 17]. The CE expands the energy as a sum of many-body configurational interactions:
A configuration represents a specific occupancy state of species its allowed sites, and describes which species sit on the -th site of the lattice. The cluster basis function is the product of site basis functions across a collection of multiple sites [57]. The average is taken over the crystal symmetry orbits , forming a complete basis to expand the energy function in the configuration space. The expansion coefficients are called effective cluster interactions (ECIs). The electrostatic energy (Ewald energy ) is included to capture long-range electrostatic interactions ( is the unscreened electrostatic energy, and is fitted as one of the ECIs).
For the simulation of atomic orderings, a cluster-expansion Hamiltonian was generated in the chemical space of Li+-Mn3+-Ti4+- O2--F-, with pair interactions up to 7.1 Å, triplet interactions up to 4.0 Å, and quadruplet interactions up to 4.0 Å based on a primitive cell of the rocksalt structure with lattice parameter Å. In total, 83 ECIs (including the constant term ) were defined, and the CE Hamiltonian was fitted with 322 unique structures. The ECIs were determined with the optimal sparseness and prediction error (with 69 non-zero elements and 8 meV/atom) with an -norm regularized regression [58]. We refer readers to Ref. [59] for details of CE in ionic systems. The CE constructions were performed using smol [60] and pymatgen for structure processing [61]. Metropolis–Hastings algorithm was used to sample the atomic configurations from the equilibrium canonical ensemble in CE-MC simulations.
Appendix C Li-vacancy phase diagram
To construct the simplified intercalation voltage profile calculations, we used an enumeration method for the generation of the delithiated structure and the construction of Li-vacancy phase diagrams. For each delithiation step as an example, it follows:
(1) Select an initial structure with composition LiTMO2, relax the structure using the product-CHGNet, and obtain the internal energy .
(2) Generate 20 different delithiated structures, each with a different vacancy occupancy and composition LiTMO2 ( may not be equal for all structures for the 0-TM to Li conversion). This generation process is referred to as an intercalation step.
(3) Relax these 20 structures using Product-CHGNet and obtain their internal energies ( denotes the index of each structure).
(4) The proposed next delithiated structure is chosen based on the minimum voltage with respect to the initial structure LiTMO2 using the following equation:
where is the internal energy of the bcc Li metal, and represents the internal energy of the (de)lithiated structures.
(5) Use the -th structure with the lowest voltage as a start and repeat Step 1 until all possible intercalation steps are consumed.
For the lithiation process, the selection criterion is changed to find the lithiated structure with the maximum voltage. It is important to note that different 0-TM channels may share the same Li atoms in an FCC framework and therefore the conversion of different 0-TM to Li thus may result in delithiated structures with varying Li concentrations in Step (2). It is more appropriate to use the minimum/maximum voltage as a selection criterion rather than the energy itself for suggesting the next step in the enumeration method.

### v2

@@
-Modeling phase transformations in Mn-rich disordered rocksalt cathodes with charge-informed machine-learning interatomic potentials
+Modeling phase transformations in Mn-rich disordered rocksalt cathodes with machine learning interatomic potentials
 Abstract
@@
 The transformed structure contains a higher concentration of non-transition metal (0-TM) face-sharing channels, which are known to improve Li transport kinetics.
-Analysis of the Mn valence distribution suggests that the appearance of tetrahedral Mn2+ is a consequence of spinel-like ordering, rather than the trigger for cation migration as previously believed.
-Calculated equilibrium intercalation voltage profiles demonstrate that the -phase, unlike the ordered spinel, exhibits solid-solution signatures during the 0-TM to Li conversion reaction.
+Analysis of the Mn valence distribution suggests that the appearance of tetrahedral Mn2+ is a consequence of spinel-like ordering, rather than the trigger for cation migration as previously suggested.
+Calculated equilibrium intercalation voltage profiles demonstrate that the -phase, unlike the ordered spinel, exhibits solid-solution signatures at low voltage.
 A higher Li capacity is obtained than in the DRX phase.
-This study provides atomic insights into solid-state phase transformation and its relation to experimental electrochemistry, highlighting the potential of charge-informed machine learning interatomic potentials for understanding complex oxide materials.
+This study provides atomic insights into solid-state phase transformation and its relation to experimental electrochemistry, highlighting the potential of machine learning interatomic potentials for understanding complex oxide materials.
 I Introduction
@@
 Disordered rocksalt (DRX) cathodes using manganese (Mn) as the earth-abundant redox-active cation have the potential to scale Li-ion energy storage to several TWh/year [4].
-Recent studies demonstrate that initiating a Mn-rich disordered rocksalt structure undergoes an in-situ transformation into a partially disordered spinel-like phase (termed as -phase) upon electrochemical cycling, resulting in increased capacity and rate capability for cathode designs [5, 6, 7, 8, 9].
+Recent studies demonstrate that Mn-rich disordered rocksalt structures undergo an in-situ transformation into a partially disordered spinel-like phase (termed the -phase) upon electrochemical cycling [5, 6, 7, 8, 9], which results in increased capacity and rate capability for cathode materials.
 This approach mitigates the typical low capacity in other DRX cathodes by promoting a new voltage feature around 4 V and an extended 3 V plateau.
-Moreover, in contrast to conventional LixMn2O4 spinel cathodes, which suffer from the two-phase reactions leading to inhomogeneity-induced stresses and mechanical degradation between , the -phase maintains a single (solid-solution) phase reaction throughout the cycling process.
+Moreover, in contrast to conventional LixMn2O4 spinel cathodes, which suffer from a two-phase reaction when fully lithiated between , leading to inhomogeneity-induced stresses and mechanical degradation, the -phase maintains a single (solid-solution) phase reaction throughout the cycling process.
 Cai et al.
-[9] demonstrates that such a series of -phase material exhibits an impressive capacity of over 280 mAh/g and an energy density exceeding 800 Wh/kg while maintaining minimal voltage decay even after extensive cycling.
-The spinel is distinct from the disordered rocksalt based on cation ordering in an FCC anion framework.
+[9] demonstrates that such -phase material exhibits an impressive capacity of over 280 mAh/g and an energy density exceeding 800 Wh/kg while maintaining minimal voltage decay even after extensive cycling.
+The spinel is distinguished from the disordered rocksalt by its cation ordering in the common FCC anion framework.
 Figure 1a–1c presents the cation orderings in a conceptualized graph representation as a projection of the 3-dimensional FCC structure shown in Fig.
-1c, including the disordered rocksalt (Fmm), spinel (Fdm), and partially disordered spinel (-phase) structures.
+1d, including the disordered rocksalt (Fmm), spinel (Fdm), and partially disordered spinel (-phase) structures.
 The green and purple spheres represent Li and Mn ions, respectively, occupying the octahedral sites (, denoted as ) of the FCC rocksalt framework.
@@
 All sites are occupied by Li, while all sites are occupied by Mn.
+Note that we refer to spinel LiMnO2 here as the product of lithiation of the traditional spinel LiMn2O4.
 The tetrahedral site neighbored by four Li is denoted as an site in a spinel.
-When the spinel is delithiated to Li0.5MnO2, the sites become vacancies and the sites are all occupied by Li.
+When the spinel is delithiated to Li0.5MnO2, the sites become vacant and the sites are occupied by Li, representing a traditional spinel structure.
 More generally, the tetrahedral site neighbored by four Li is also termed a 0-TM channel.
 Urban et al.
-[10] demonstrated that long-range connected 0-TM percolating channels are essential for facile Li diffusion in an FCC framework, with the spinel ordering having the highest amount of 0-TM channels in theory.
+[10] demonstrated that long-range connected 0-TM percolating channels are essential for facile Li diffusion in an FCC framework, with the spinel ordering having the highest amount of 0-TM channels of any ordering.
 The partially disordered spinel is conceptually illustrated in Fig.
-1c as an intermediate state between the spinel and disordered ordering.
-The phase transformation from DRX to -phase is related to the transition metal (TM) migration of the mobile Mn ion when neighboring cation vacancies are created [11].
-The TM can take the pathway connected by the octahedral-tetrahedral-octahedral (--) sites in an FCC framework.
-For the migration process, it is typically concluded that Mn migration is strongly coupled with its charge state, Mn4+ is generally considered immobile, while Mn3+ and Mn2+ are considered more mobile according to the nudged-elastic band calculations [12, 13, 14].
+1c as an intermediate state between the spinel and the disordered rocksalt.
+The phase transformation from the DRX to -phase is related to the transition metal (TM) migration of the mobile Mn ion when neighboring cation vacancies are created [11].
+The TM migration can occur along the pathway connected by the octahedral-tetrahedral-octahedral (--) sites in an FCC framework.
+The Mn migration barrier is strongly coupled with its charge state: Mn4+ is generally considered immobile, while Mn3+ and Mn2+ are considered more mobile according to the nudged-elastic band calculations [12, 13, 14].
 However, simulating these charge-coupled dynamics has long been challenging, as the required time and length scales are inaccessible to conventional ab initio methods.
-Previously, cluster expansion (CE) methods have been shown successfully as a surrogate model of density functional theory (DFT) calculations for modeling the configurational thermodynamics in complex oxides via Monte Carlo (MC) simulation at scale [15, 16, 17].
+Previously, cluster expansion (CE) methods have successfully acted as a surrogate model to density functional theory (DFT) calculations for modeling the configurational thermodynamics in complex oxides via Monte Carlo (MC) simulation at scale [15, 16, 17].
 Nonetheless, the limited expressibility and coarse-grained nature of these models render them less adequate for capturing the complex charge-coupled dynamical processes involved in TM migrations [18, 9].
@@
 It identified the timescale discrepancies between charge disproportionation and ion hops, suggesting that the emergence of Mn2+ is correlated to long-range spinel-like ordering but not solely to the local hopping into the tetrahedral occupancy.
-In this study, we extend the computational study of the phase transformation in LixMn0.8Ti0.1O1.9F0.1 (LMTOF), which is a Mn-rich DRX cathode.
+In this study, we study the complex phase transformation in LixMn0.8Ti0.1O1.9F0.1 (LMTOF), which is a Mn-rich DRX cathode material.
 We parametrized a CHGNet-MLIP using r2SCAN-DFT calculations for various atomic configurations to sample the potential energy surface in this chemical space.
 The fine-tuned CHGNet-MLIP was used to implement charge-informed MD simulations starting from a disordered phase in Li0.6Mn0.8Ti0.1O1.9F0.1.
-The MD simulations revealed a phase transformation from a disordered rocksalt to a spinel-like () phase.
-From the magnetic moment predictions, we demonstrate that the Mn migration is mainly contributed by Mn3+ at the beginning, while Mn2+ collectively emerges as the spinel-like ordering forms.
-Furthermore, we investigated the impact of this partial disorder on the electrochemical intercalation voltage profiles based on the structures obtained from the MD simulations.
+The MD simulations reveal a phase transformation from a disordered rocksalt to a spinel-like () phase.
+From the magnetic moment predictions, we demonstrate that the Mn migration mainly occurs as Mn3+ in the initial part of the transformation, while Mn2+ collectively emerges as the spinel-like ordering forms.
+Furthermore, we investigated the impact of this partial disorder on the electrochemical intercalation voltage profiles using the structures obtained from the MD simulations.
 Our work provides insights into the application of MLIPs for studying complex energy materials and offers a mechanistic understanding of the partial cation-disorder effect on intercalation electrochemistry at the atomic scale.
@@
 After message-passing layers, the node features are linearly projected onto magnetic moments by , and the total energy is calculated as a sum of non-linear node projections by over all atoms with an additional message-passing layer
-The nodes are regularized by DFT magnetic moments to carry rich information about both local ionic environments and to serve as a proxy for the atomic charge.
+The nodes are regularized by DFT magnetic moments to carry rich information about both local ionic environments.
+The valence state of TM is inferred from the predicted magnetic moment as a proxy of the atomic charge, which is a direct output of the MLIP rather than an input parameter.
 The forces and stress are calculated via auto-differentiation of the total energy with respect to the atomic Cartesian coordinates and strain:
-This approach allows CHGNet to capture the complex relationships between charge distributions and energetics with given atomic configurations.
+This approach allows CHGNet to capture the complex relationships between charge distributions and energetics with given atomic configurations, which is essential to properly model the kinetic path of evolving TM oxides [23].
 II.2 Potential energy surface sampling
-While CHGNet [21] as a pre-trained foundation potential achieves reasonable performance in predicting materials stability on out-of-distribution datasets [23], recent benchmarks observed a systematic softening of the ML predicted potential energy surface (PES) compared to the DFT ground truth, particularly for atomic configurations from MD simulations with high energy [24, 25].
+CHGNet [21] as a pre-trained foundation potential achieves reasonable performance in predicting materials stability on out-of-distribution datasets [24].
+However, recent benchmarks observed a systematic softening of the ML predicted potential energy surface (PES) compared to the DFT ground truth, particularly for atomic configurations with high energy [25, 26].
 To improve simulation accuracy, we performed multi-faceted sampling of the PES for the Li–Mn–Ti–O–F chemical space [27], as follows:
 Exisiting dataset – We collected the compounds in the Materials Project database within the Li–Mn–Ti–O–F chemical space that have decomposition energy (energy above the convex hull, ) lower than 0.1 eV/atom as a general coverage of such chemical space [28].
-Furthermore, we adopted the structures from the training dataset of previously reported cluster expansion (CE) models for DRXs containing Mn2+/Mn3+ as redox-active species (DRX-CE) [28, 29].
+Furthermore, we adopted the structures from the training dataset of previously reported cluster expansion (CE) models for disordered rocksalts containing Mn2+/Mn3+ as redox-active species (DRX-CE) [29, 30].
 All these structures in the DRX-CE dataset have rocksalt lattices but different cation and anion orderings.
@@
 Partially disordered structures – Given that existing datasets do not include atomic configurations specifically representing partially disordered structures with spinel-like orderings, we developed a custom sampling procedure to generate these configurations.
-(1) Monte Carlo (MC) simulations were employed to obtain the disordered rocksalt atomic configurations equilibrate at finite temperatures, which were generated in supercells of the primitive spinel cell (see Appendix B for details).
-Given these disordered structures, we compared the occupancy of the spinel supercell and the disordered rocksalt in such a supercell lattice.
-The octahedral cation sites are categorized into and sites.
-Under this classification, a random distribution of TM yields the disordered structure, while an increased number of TM occupations on sites indicates more spinel-like ordering.
-To generate different degrees of spinel order, we applied site exchanges between the two nearest octahedral sites to transform the disordered structure towards the spinel ordering through either (a) an exchange between a TM and a Li via a Li3M tetrahedron; or (b) an exchange between a TM and a Li.
-The (a) step is to create more spinel-like ordering and the (b) step is to exchange the cation ordering within the sites to potentially create more available Li3M tetrahedron.
-We assume that TM migration occurs through the -- path with no other TM neighbors in the delithiated states (i.e., through the Li3M tetrahedron, Fig.
-3).
-By iteratively applying site exchanges within different Li3M tetrahedra, one can obtain a collection of structures ranging from 0% (disordered) to nearly 100% spinel ordering.
-TM migration pathways – Since TM migration is a rare, high-energy-barrier event [9], especially for unfavorable valence states like Mn4+, we implemented a passive sampling approach to effectively model the diffusion kinetics with the following steps:
-For each site exchange step via the Li3M tetrahedron, three different states are generated as A (initial state, TM occupied on site), B (intermediate state, TM occupied on site), and C (final state, TM occupied on site), which are shown in Fig.
+(a) Monte Carlo (MC) simulations were employed to obtain disordered rocksalt atomic configurations in a supercell of the primitive spinel cell by equilibration at high temperatures (see Appendix B).
+(b) The octahedral cation sites are then categorized into and sites by overlaying the template of the perfectly ordered spinel.
+Under this classification, a random distribution of TM yields the disordered structure, while preferential TM occupation of sites indicates more spinel-like ordering.
+To generate different degrees of spinel order, we applied site exchanges between two nearest neighbor octahedral sites to transform the disordered structure towards the spinel ordering through: (a) An exchange between a TM and a Li in a Li3M tetrahedron.
+We selected a Li3M tetrahedron that contains three Li atoms (including at least one Li) and one metal atom.
+A site exchange is applied to the Li and the TM.
+(b) An exchange between a TM and a Li.
+Step (a) increases the spinel-like order, but it may halt if no suitable Li3M tetrahedra are available to facilitate the exchange.
+To address this, Step (b) rearranges cations within the sites.
+While this internal exchange does not change the overall degree of spinel ordering, its purpose is to create new Li3M tetrahedra, thereby enabling further transformations via Step (a).
+By iteratively applying site exchanges within different Li3M tetrahedra, one can obtain a collection of structures ranging from disordered to spinel ordering.
+TM migration pathways – TM migration is a rare, high-energy-barrier event [9], especially for unfavorable valence states such as Mn4+.
+To effectively model the diffusion kinetics, we implemented a passive sampling approach with the following steps:
+A TM can only migrate from its octahedral site into or through a tetrahedral site which has no other face-sharing TMs.
+In the fully lithiated state, there are Li3M tetrahedra where the tetrahedron is surrounded by only one TM and 3 Li.
+For each migration event via a Li3M tetrahedron, three different states are generated as A (initial state, TM occupied on site), B (intermediate state, TM occupied on site), and C (final state, TM occupied on site), which are shown in Fig.
 3.
-The other neighboring Li atoms are removed to create vacancies, for which we assume no face-sharing between the migrating TM and cations.
-The structures in the A/B/C states were relaxed using the Trial-CHGNet fitted with structures from the existing datasets.
+The Li atoms face-sharing with the site are removed to create vacancies so that there are no face-sharing ions with the migrating TM and cations.
+The structures in the A/B/C states were relaxed using the Trial-CHGNet.
 After structure relaxation, five images interpolated along A-B (- path) and B-C (- path) are evaluated with static DFT calculations, analogous to the intermediate states for nudged elastic band (NEB) calculations.
-Furthermore, we follow the same approach to generate the direct migration from the - path (A-C) for labeling high-energy and unfavorable states.
+Furthermore, we follow the same approach to generate states along a direct migration between sites, which cuts through their shared edge (A-C) so that high-energy states are considered in the MLIP fitting [31].
 In addition, we considered two different delithiation levels for sampling TM migration pathways: (a) A dilute level, where only three Li ions are removed to create a diffusion channel for TM migration (e.g., a pristine supercell structure with Li44Mn32Ti4O76F4 yields a delithiated structure with Li41Mn32Ti4O76F4).
-This level allows for the investigation of local TM hopping energetics.
+This level allows for the investigation of the environment dependence of the TM hopping energetics.
 (b) A medium delithiation level, where 0.5 Li per formula unit (f.u.) is removed from the structure to create delithiated states (e.g., Li ions are randomly removed to create a structure with Li24Mn32Ti4O76F4).
 This level allows the sampling of delithiated states, representing the charging process in a battery.
-By sampling partially disordered structures at these two delithiation levels, we include a more comprehensive description of the cation ordering energetics, TM hop, and relevant local environments in which TM hopping occurs at the composition that is energetically favorable for spinel formation (i.e., the composition close to Li0.5TMO2).
+By sampling partially disordered structures at these two delithiation levels, we include a more comprehensive description of the cation ordering energetics, TM hop, and relevant local environments in which TM hopping occurs at the composition that is energetically favorable for spinel formation (i.e., the composition close to Li0.5TMO2) [32].
 By adding the static DFT calculations of these sampled atomic configurations to the relaxation trajectories, another MLIP was parameterized (denoted as Sample-CHGNet) to perform MD simulations for enhanced sampling.
-MD trajectories – To further sample the thermally excited atomic configurations at finite temperatures, we employed Sample-CHGNet to perform molecular dynamics (MD) simulations with disordered structures across various compositions.
+MD trajectories – To further sample the thermally excited atomic configurations at finite temperatures, we employed Sample-CHGNet to perform molecular dynamics (MD) simulations on disordered structures across various compositions.
 These structures were first generated in the pristine (lithiated) state from the CE-MC simulations at K using supercells of and formula units (f.u.), including Li1.2Mn0.5Ti0.3O1.9F0.1, Li1.2Mn0.6Ti0.2O1.8F0.2, Li1.1Mn0.8Ti0.1O1.9F0.1, and Li1.1Mn0.7Ti0.2O2.0.
@@
 III Results
-To investigate the structural evolution, we performed charge-informed molecular dynamics from the disordered structure to monitor the phase change to the -phase.
+To investigate the structural evolution as a disordered rocksalt transforms, we performed charge-informed molecular dynamics from the disordered structure to monitor the phase change to the -phase.
 The initial (disordered) structure was generated by sampling from the equilibrium ensemble of CE-MC simulations at K.
-The MC structure represents the pristine (lithiated) composition (Li1.1Mn0.6Ti0.1O1.9F0.1) with a supercell containing 320 atoms.
-Based on the pristine structure, a delithiated structure (LiMn0.6Ti0.1O1.9F0.1, where represents cation vacancy) was then obtained by randomly removing Li ions and relaxing the delithiated configuration using the Product-CHGNet with a force convergence criterion of eV/Å.
+The MC structure represents the pristine (lithiated) composition (Li1.1Mn0.8Ti0.1O1.9F0.1) with a supercell containing 320 atoms.
+Based on the pristine structure, a delithiated structure (LiMn0.8Ti0.1O1.9F0.1, where represents a cation vacancy) was then obtained by randomly removing Li ions and relaxing the delithiated configuration using the Product-CHGNet with a force convergence criterion of eV/Å.
 After relaxation, MD simulations were subsequently performed for ns with a time step of fs under the NVT ensemble at K.
@@
 Figure 4a presents the simulated X-ray diffraction (XRD) patterns derived from the structures within the time interval of to ns.
-To analyze the evolution of cation ordering on the FCC lattice, the atomic positions from the MD snapshots were refined to their nearest ideal Wyckoff sites.
+To analyze the evolution of cation ordering on the FCC lattice, the atoms in the MD snapshots were moved to their nearest ideal Wyckoff sites.
 The gray dashed lines in Fig.
@@
 At the onset of the MD simulation ( ns), there are no obvious peaks indicating spinel ordering.
-After approximately ns, the intensity of the spinel peaks increases and becomes more significant in subsequent time scales as the MD simulation proceeds, suggesting the formation of the spinel-like -phase.
+After approximately ns, the intensity of the spinel peaks increases and becomes more significant as the MD simulation proceeds, suggesting the formation of the spinel-like -phase.
 Figure 4b illustrates the energy landscape across the simulated MD trajectories.
-The black points denote the potential energies, while the colored points represent the total energies.
+The black points denote the potential energy, while the colored points represent the total energy.
 The energies are depicted relative to their values at the initial state.
-A notable decrease in energy (approximately 60 meV/anion) is observed for the -phase following a ns MD simulation.
+A notable decrease in energy (approximately 60 meV/anion) is observed after ns when the -phase has formed.
 This energy reduction indicates that spinel-like cation ordering is thermodynamically favored at the composition Li0.6Mn0.8Ti0.1O1.9F0.1 when sufficient cation vacancies are present.
@@
 To study the change in local cation ordering, we analyzed the short-range order (SRO) indicated by the tetrahedron occupancy as a function of MD simulation time.
-Figure 4c presents the occurrence of 0-TM (solid green line) and 1-TM tetrahedral (solid purple line) channels in the MD structure, as a ratio to all interstitial tetrahedral sites.
-The 0-TM channel has a low migration barrier that kinetically favors fast Li-ion transport [31].
+Figure 4c presents the occurrence of 0-TM (solid green line) tetrahedra in the MD structure, as a ratio to all tetrahedral sites.
 To compare with disordered phases, the dashed green lines in Fig.
-4c indicate the random limit of the 0-TM channel for two Li contents ( for Li1.0 and for Li1.1 at the fully lithiated level) for reference.
-At the beginning of the MD simulation ( ns), the occurrence of 0-TM channels is approximately 0.03.
-This value is significantly lower than that of the random limit of Li1.1TMO2 (0.09), where the value is estimated based on a rocksalt structure with 0.1 excess Li and fully random cation ordering.
-The low value of 0-TM content in the initial structure indicates reduced percolating Li content and suggests that SRO is unfavorable for long-range Li percolation, which is consistent with previous predictions in Mn-Ti-based DRXs [28, 29].
-As the phase transformation progresses, the occurrence of 0-TM channels increases, reaching approximately 0.06 after ns of MD simulation.
-This rise in the number of 0-TM channels agrees with the high 0-TM content in the spinel structure, indicating the emergence of spinel-like ordering.
+4c indicate the random limit of the 0-TM tetrahedra for two Li contents ( for Li1.0 and for Li1.1 at the fully lithiated level) for reference.
+At the beginning of the MD simulation ( ns), the occurrence of 0-TM tetrahedra is approximately 0.03.
+This value is significantly lower than that of the random limit for Li1.1TMO2 (0.09).
+The low value of 0-TM content in the initial structure indicates that SRO is unfavorable for long-range Li percolation, which is consistent with previous predictions in Mn-Ti-based DRXs [29, 30].
+As the phase transformation progresses, the occurrence of 0-TM tetrahedra increases, reaching approximately 0.06 after ns of MD simulation.
+This rise in the number of 0-TM tetrahedra agrees with the high 0-TM content in the spinel structure, indicating the emergence of spinel-like ordering.
 III.2 Charge distribution with ion rearrangement
-To understand the correlation between Mn migration and its heterovalent state [9, 32], we inferred atomic charge states from the magnetic moments predicted by CHGNet.
+To understand the correlation between Mn migration and its valence state [9, 34], we inferred atomic charge states from the magnetic moments predicted by CHGNet.
 Figure 5a presents the site-wise distribution of magnetic moments for Mn ions.
-The heterovalent states of Mn are categorized as follows: Mn, Mn, and Mn.
-At the beginning of the MD simulation, Mn predominantly exists in valence states of +3 and +4.
+The valence states of Mn are categorized as follows: Mn, Mn, and Mn.
+At the beginning of the MD simulation, Mn predominantly exists in valence states of +3 and +4, consistent with the partial delithiation of the material.
 As the phase transformation progresses, the concentration of Mn2+ gradually increases, as shown by the increased frequency of the highlighted circles in Fig.
@@
 is plotted in Fig.
-5b as an indication of hopping events to tetrahedral sites.
-Although previous nudged elastic band calculations reveal that Mn2+ has a lower migration barrier [11, 9], the atomic charges predicted by CHGNet show that most of the tetrahedral Mn are Mn from to ns during the MD simulation.
+5b.
+Although previous NEB calculations reveal that Mn2+ has the lower migration barrier of all Mn valence states [11, 9], the atomic charges predicted by CHGNet show that most of the tetrahedral Mn are Mn from to ns during the MD simulation.
 This suggests that Mn3+ may migrate without the need for charge disproportionation () as proposed by Reed et al.
@@
 This evidence suggests that Mn3+ migration initiates the transformation before long-range spinel-like order is established.
-The subsequent emergence of tetrahedral Mn2+ appears to be a consequence of this long-range ordering (ns-scale) rather than a prerequisite for local cation hops (ps-scale).
+The subsequent emergence of tetrahedral Mn2+ appears to be a consequence of this long-range ordering rather than a prerequisite for local cation hops.
 The simulation results are consistent with previous MD simulations of the LixMnO2 phase transformation revealed by Deng et al.
@@
 III.3 Intercalation voltage profile of -phase
-To evaluate the effect of partial disorder on the electrochemical performance of the transformed structure from MD, we computed the intercalation voltage profiles using Product-CHGNet.
-We focused on two characteristic reactions in an ordered spinel LiMnO: (1) a 0-TM conversion reaction at a lower voltage, where the occupancies of Li ions are transformed from the octahedral sites to the tetrahedral sites collectively:
-and (2) a conversion at a higher voltage, where Li is intercalated on the site.
-It is important to note that Eq.
-(3) is defined for an independent 0-TM channel (i.e., one that does not share Li with another tetrahedral channel).
-In an ordered LiMnO2 spinel, each Lioct is face-sharing with two 0-TM channels, where removing one Li from such a 0-TM channel results in a high defect energy [18].
-The delithiation is, therefore, a collective conversion that exhibits a two-phase reaction between Li0.5MnO2 and LiMnO2 with a characteristic 3 V voltage plateau [33].
-Using Eq.
-(3), we enumerated different Li-vacancy configurations with fixed TMs and an anion framework (topotactically) at each (de)lithiation level based on the MD structure.
-The Product-CHGNet was used to perform structure relaxation for both lithiation and delithiation paths (see Appendix C for details).
+To evaluate the effect of partial disorder on electrochemical performance, we computed the intercalation voltage profiles of the transformed -phase structure using Product-CHGNet.
+Specifically, we focus on the voltage characteristic of ordered spinel LixMn2O4 at lower voltage (for ), in which the occupancies of Li ions are transformed collectively from the tetrahedral sites to the empty octahedral sites upon lithiation.
+The reason this happens collectively in a well-ordered spinel is that (tet) and (oct) sites cannot be simultaneously occupied.
+Hence, upon lithiation the system empties the sites and occupies .
+However, in a (partially) disordered rocksalt structure, the , , and crystallographic labels cannot be defined precisely.
+For this reason, we define a “0-TM” site as any interstitial tetrahedral site that is exclusively face-shared by four non-transition metal cations (i.e., Li or vacancy).
+This definition captures the essential local environment of the site in an ordered spinel which also has no transition metal neighbors.
+It also generalizes to disordered rocksalts where the connectivity of 0-TM sites is essential for evaluating the percolating Li content [10].
+Using this more general description, we can expect two key intercalation processes:
+(a) A lower-voltage reaction whereby a tetrahedral Li is removed and its octahedral sites become filled with Li. For such an isolated 0-TM site, this reaction can be written as:
+(b) A higher-voltage reaction corresponding to the direct insertion of Li into an empty 0-TM tetrahedral site:
+This framework allows us to consistently compare the electrochemical behavior of the ordered spinel, disordered rocksalt, and the intermediate -phase.
+We generated a series of Li-vacancy configurations to simulate the electrochemical (de)lithiation process topotactically.
+Starting with the fully lithiated structure obtained from MD, we held the TM and anion sublattices fixed.
+At each (de)lithiation level, distinct Li-vacancy orderings were created by selecting different 0-TM sites within the host structure and applying the local transformation described in Eq.
+(3) and (4).
+The Product-CHGNet was used to perform structure relaxation for both lithiation and delithiation paths (see Appendix C).
 The energies of these (de)lithiated structures were used to construct the convex hull and the corresponding voltage profile as shown in Fig.
 6a and Fig.
-6b (see Methods).
+6b.
 For the region with high Li concentration (), the 0-TM to conversion exhibits nearly solid-solution behavior, as evidenced by the convex hull being connected by several ground states in Fig.
 6a.
-Particularly, for the early stage of delithiation (), the voltage exhibits a flat plateau with a minor slope, indicating that the delithiation through Eq.
+Particularly, for the early stage of delithiation (), the voltage exhibits a pseudo plateau with minor slope, indicating that the delithiation through Eq.
 (3) occurs in similar local chemical environments and exhibits similar chemical potentials (voltages).
-We found that the voltage at is around 2.40 V, which is slightly lower than the experimental value.
+The computed voltage at is around 2.40 V, which is slightly lower than the experimental value.
 This discrepancy is likely attributable to the residual self-interaction error common in DFT functionals, even for r2SCAN, when a Hubbard U correction is not applied [35, 36].
 Although the SCAN-based functional removes more self-interaction than previous LDA and GGA functionals [37], it does not fully remove it and underestimates the intercalation voltage, which is consistent with previous reports [18, 17].
-Upon reaching , lithium occupies all tetrahedral 0-TM channels (Li) in addition to the octahedral Li not in the 0-TM channels.
+Upon reaching , lithium occupies all tetrahedral 0-TM sites (Li) in addition to the octahedral Li not in the 0-TM sites.
 Two competing delithiation processes are available, involving either the removal of Li or Li.
-We computed both pathways and found that the low-energy configurations near the convex hull within the range correspond to the removal of Li, while the high-energy configurations (represented by dark red markers in Fig.
+We computed both pathways and found that the low-energy configurations near the convex hull in the range correspond to the removal of Li, while the high-energy configurations (represented by dark red markers in Fig.
 6a) correspond to the removal of Li.
 These results indicate that even in a structure with partial spinel ordering, Li has a preference for delithiation before Li.
-In contrast, the delithiation in spinel LixMnO2 for entirely corresponds to the removal/insertion of Li with a V plateau.
+In contrast, the delithiation in ordered spinel LixMnO2 for entirely corresponds to the removal/insertion of Li with a V plateau.
 We also observe a voltage increase of V around (the end of 0-TM to Li conversion).
-The magnitude is not as large as that in spinel LiMnO2 ( V) or in the spinel-ordered LixMn0.8Ti0.1O1.9F0.1 (see Fig.
-7c), suggesting the absence of two-phase reactions in the -phase and less inhomogeneity that could lead to battery performance degradation.
-To further demonstrate the effect of partial disorder, we evaluated the intercalation voltage profiles of Li1.1Mn0.8Ti0.1O1.9F0.1 in both the DRX and spinel phases for comparison.
-The DRX structure was selected from the structure derived from the CE-MC simulations (i.e., the configuration at s in the MD simulation).
-The spinel Li1.1Mn0.8Ti0.1O1.9F0.1 was generated through another CE-MC simulation without exchanges between the and sites, where all sites are occupied by Li1.0 and sites are occupied by Mn0.8, Ti0.1, and Li0.1.
+The magnitude is not as large as that in spinel LiMn2O4 ( V) or in the spinel-ordered LixMn0.8Ti0.1O1.9F0.1 (see Fig.
+7d), suggesting the absence of two-phase reactions in the -phase.
+To further demonstrate the effect of partial disorder, we evaluated the intercalation voltage profiles of Li1.1Mn0.8Ti0.1O1.9F0.1 in both the DRX and spinel phases as a comparison.
+The DRX structure was selected from the CE-MC simulations (i.e., the configuration at s in the MD simulation).
+The spinel Li1.1Mn0.8Ti0.1O1.9F0.1 was generated through a separate CE-MC simulation without exchanges between the and sites, so that all sites are occupied by Li1.0 and sites are occupied by Mn0.8, Ti0.1, and Li0.1, similar to a Li-excess lithiated spinel.
 The MC simulation allows site exchange within sites and anion sites, which is designed to capture the underlying SRO of anions and cations in DRXs (e.g., Li-F SRO) [38].
-The intercalation at high Li content in both DRX and spinel phases only included the 0-TM conversion reaction as described in Eq.
+The intercalation at high Li content in both DRX and spinel phases only included the 0-TM site conversion reaction as described in Eq.
 (3).
 Figure 7a and 7b illustrate the formation energy convex hull and the voltage profile of the DRX phase, with the -phase plotted as a red line for comparison.
-Due to the insufficient number of 0-TM channels in the DRX structure, a relatively limited Li capacity from the 0-TM conversion is achieved (approximately 0.3 per f.u., as denoted by the green dashed line in Fig.
+As there are few 0-TM sites in the DRX structure, a relatively limited Li capacity from the 0-TM conversion is achieved (approximately 0.3 per f.u., as denoted by the vertical green dashed line in Fig.
 7a and 7b.
-The absence of spinel-like ordering results in a profile marked by a mildly steep increase as delithiation progresses.
-Notably, the overall voltage profile between of the DRX phase is higher than that of the -phase, suggesting a restricted amount of accessible capacity within an operational voltage window, particularly the capacity around V [9, 8].
-In contrast, Figure 7b depicts the calculated results for spinel LixMn0.8Ti0.1O1.9F0.1.
-The voltage profile clearly displays a two-phase reaction between Li1.1 and Li0.6 with a linear convex hull in Fig.
+The absence of spinel-like ordering results in a profile marked by a mildly steep voltage increase as delithiation progresses.
+Notably, the overall voltage profile of the DRX phase between is higher than that of the -phase, suggesting a restricted amount of accessible capacity around V [9, 8].
+In contrast, Figure 7d depicts the calculated results for spinel LixMn0.8Ti0.1O1.9F0.1.
+The voltage profile clearly displays a two-phase reaction between Li1.1 and Li0.6 consistent with the linear convex hull in Fig.
 7c.
-Due to the absence of disorder, the spinel phase exhibits a lower average voltage with a two-phase reaction between , with a voltage step more than 1 V presented at the end of 0-TM conversion.
+Due to the absence of disorder, the spinel phase exhibits a lower average voltage with a two-phase reaction between , and a voltage step larger than 1 V at the end of 0-TM conversion.
 For , the intercalation corresponds to Li, which shows a similar high-voltage feature as the -phase.
-The effect of cation disorder on the spinel structure demonstrated in our simulations is in full agreement with the previous findings on the cation disorder effects in LiMnO2 spinel by Chen et al.
+The effect of cation disorder on the spinel structure demonstrated in our simulations is in agreement with the previous findings on the cation disorder effects in LiMnO2 spinel by Chen et al.
 [18].
 IV Discussion
-In the evolving landscape of modern battery materials, the complexity of components and configurations exemplified by DRX cathode materials necessitates advanced computational models.
-Particularly, Mn-rich DRXs exhibit a phase transformation to a partially disordered spinel phase (-phase) via in situ electrochemical cycling [9, 8, 38] or chemical delithiation [39].
-This transformation is challenging for conventional modeling approaches to capture, particularly the kinetic pathways of TM migration in such a complex oxide.
+In the evolving landscape of modern battery materials, the multirole of components and complexity of configurations exemplified by DRX cathode materials necessitates advanced computational models.
+Particularly, Mn-rich DRXs exhibit a phase transformation to a partially disordered spinel phase (-phase) via in situ electrochemical cycling [9, 8, 39] or chemical delithiation [40, 41].
+This transformation is challenging for conventional modeling approaches, as it requires capturing the kinetic pathway of TM migration in the disordered environment.
 While ab initio molecular dynamics could theoretically address this issue, its computational cost renders it impractical for probing large systems or long timescales.
 Alternatively, kinetic Monte Carlo (kMC) methods with CE models have been proposed to model cation ordering changes in both cathodes and solid electrolytes [42, 43].
-However, parameterizing CEs for disordered oxides/oxyfluorides and accurately calculating migration barriers in diverse local environments remains a non-trivial endeavor [42].
-To address these challenges, we fine-tuned a charge-informed foundation MLIP with sampled r2SCAN-DFT calculated energies, forces, stresses, and magnetic moments.
-The multi-step PES sampling captures the complex interactions from diverse chemical environments and heterovalent states of TMs.
+However, parameterizing CEs for disordered oxides/oxyfluorides and accurately calculating migration barriers in diverse local environments with DFT remains a non-trivial endeavor [44].
+To address these challenges, we fine-tuned a foundation MLIP with sampled r2SCAN-DFT calculated energies, forces, stresses, and magnetic moments.
+The multi-step PES sampling captures the complex interactions from diverse chemical environments and different valence states of TMs.
 Our MD simulations using the fine-tuned CHGNet successfully captured the phase transition from a disordered to a spinel-like configuration in Li0.6Mn0.8Ti0.1O1.9F0.1.
-The simulations reveal that the formation of the -phase promotes the occurrence of 0-TM channels, which play a crucial role in facilitating long-range Li-ion transport, whereas the DRX phase exhibits limited availability of these channels.
-This result agrees with the improved high-rate performance observed in experiments [9, 8, 39].
-By analyzing the electrochemical intercalation voltage profiles of the -phase and comparing them to those of the DRX and spinel phases, two conclusions can be drawn from the simulation results: (1) The presence of partial cation disorder in the -phase effectively eliminates the two-phase reaction, as evidenced by the absence of a sharp voltage increase after Li0.5 removal (see Fig.
-7c).
-(2) The introduction of spinel-like ordering increases the 0-TM occurrence (as indicated by the increased number of 0-TM channels in Fig.
+The simulations reveal that the formation of the -phase from DRX promotes the increase of 0-TM configurations on tetrahedra, which play a crucial role in facilitating long-range Li-ion transport.
+This result is consistent with the improved high-rate performance observed in experiments [9, 8, 40].
+By analyzing the simulated electrochemical intercalation voltage profiles of the -phase and comparing them to those of the DRX and spinel phases, two conclusions can be drawn: (1) The presence of partial cation disorder in the -phase effectively eliminates the two-phase reaction, as evidenced by the absence of a sharp voltage increase after Li0.5 removal (see Fig.
+7d).
+This is consistent with earlier predictions on partially disordered spinels [18].
+(2) The introduction of spinel-like ordering increases the 0-TM occurrence (as indicated by the increased number of 0-TM tetrahedra in Fig.
 4c) and thus expands the available Li capacity from the 0-TM to Li conversion reaction around 3 V (denoted by the green dashed line in Fig.
 6b).
-The Li capacity from 0-TM to Li conversion is even higher than that of the spinel in Fig.
-5c since the 0-TM channels in partially disordered phases are less correlated with each other compared to the spinel, where each Li is shared by two 0-TM channels.
+The Li capacity from 0-TM to Li conversion in the -phase is even higher than that of the spinel in Fig.
+7d since the 0-TM sites in partially disordered phases are less spatially correlated with each other compared to the spinel, where each Li is shared by two 0-TM sites.
 Although this study represents an integrated investigation of phase transformation in complex oxides combining MC simulations for configurational thermodynamics and MLIP-MD simulations for kinetic modeling, several limitations were identified that could be further addressed in future studies.
@@
 Achieving ab initio accuracy while enabling simulations beyond typical MD timescales (such as using generative models for dynamic trajectories) holds promise for investigating these processes [47, 48, 49].
-Regarding model fine-tuning, recent advances such as active learning with Bayesian force fields [48], stratified sampling from MLIP-driven MD [49], and diffusion-based generative models for mapping reaction pathways [50] provide more efficient sampling techniques beyond the passive sampling implemented in this work, which could enhance MLIP training accuracy by exploring the PES more effectively.
-In summary, we demonstrated the application of charge-informed MLIP to model the complex phase transformation from DRX to spinel-like -phase in Li-Mn-Ti-O-F systems.
+Regarding model fine-tuning, recent advances such as active learning with Bayesian force fields [50], stratified sampling from MLIP-driven MD [51], and diffusion-based generative models for mapping reaction pathways [52, 53] provide more efficient sampling techniques beyond the passive sampling implemented in this work, which could enhance MLIP training accuracy by exploring the PES more effectively.
+Finally, it is important to clarify that the term ”charge” refers to the valence state (nominal charge) of the TM, which is inferred from the predicted magnetic moments.
+The current model does not explicitly compute long-range electrostatic (Coulombic) interactions.
+Looking forward, the model’s capabilities could be significantly enhanced by incorporating response charge prediction, such as Latent Ewald Summation (LES) [54].
+LES integrates long-range electrostatics and the prediction of Born effective charges [55], which has demonstrated a universal augmentation for mainstream MLIPs [56].
+We believe an LES-augmented CHGNet would yield a more physically robust description of charge, thereby improving the modeling of complex phenomena such as cation migration during phase transformations in oxides.
+In summary, we demonstrated the application of a MLIP to model the complex phase transformation from DRX to spinel-like -phase in Li-Mn-Ti-O-F systems.
 The MLIP-MD simulations revealed the mechanism of spinel-like ordering formation and its effect on electrochemical properties.
-As computational capabilities and new methodologies emerge–particularly in accelerating MLIP simulations and in generative modeling–these advances will pave the way for more accurate and efficient modeling of phase transformations in next-generation energy storage materials.
+As computational capabilities and new methodologies emerge–particularly in accelerating MLIP simulations and in generative modeling–these advances will pave the way for more accurate and efficient modeling of phase transformations in new energy storage materials.
 V Acknowledgments
@@
 DE-AC0205CH11231.
-The computational modeling in this work was supported by the computational resources provided by the Extreme Science and Engineering Discovery Environment (XSEDE), supported by National Science Foundation grant number ACI1053575; the National Energy Research Scientific Computing Center (NERSC); and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory.
+The computational modeling in this work was supported by the computational resources provided by the Extreme Science and Engineering Discovery Environment (XSEDE), supported by National Science Foundation grant number ACI1053575; the National Energy Research Scientific Computing Center (NERSC); the National Renewable Energy Laboratory (NREL) clusters under the drx allocation, and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory.
 The authors thank Zijian Cai for the valuable discussions.
+Data Availability
+The supporting data and code are available at https://github.com/zhongpc/drx2delta.
 Appendix A DFT Calculations
-All DFT calculations were performed with the VASP package using the projector-augmented wave method [51, 52], a plane-wave basis set with an energy cutoff of 680 eV, and a reciprocal space discretization of 25 k-points per Å-1.
+All DFT calculations were performed with the VASP package using the projector-augmented wave method [57, 58], a plane-wave basis set with an energy cutoff of 680 eV, and a k-point sampling grid spacing of 0.25 Å-1.
 The calculations were converged to eV in total energy for electronic loops and 0.02 eV/Å in interatomic forces for ionic loops.
@@
 Appendix C Li-vacancy phase diagram
-To construct the simplified intercalation voltage profile calculations, we used an enumeration method for the generation of the delithiated structure and the construction of Li-vacancy phase diagrams.
-For each delithiation step as an example, it follows:
-(1) Select an initial structure with composition LiTMO2, relax the structure using the product-CHGNet, and obtain the internal energy .
+To construct the simplified intercalation voltage profiles, we used an enumeration method for the generation of the delithiated structure and the construction of Li-vacancy phase diagrams.
+For each delithiation level, the following steps are performed:
+(1) Select an initial structure with composition LiTMO2, relax the structure using Product-CHGNet, and obtain the internal energy .
 (2) Generate 20 different delithiated structures, each with a different vacancy occupancy and composition LiTMO2 ( may not be equal for all structures for the 0-TM to Li conversion).
@@
 For the lithiation process, the selection criterion is changed to find the lithiated structure with the maximum voltage.
-It is important to note that different 0-TM channels may share the same Li atoms in an FCC framework and therefore the conversion of different 0-TM to Li thus may result in delithiated structures with varying Li concentrations in Step (2).
+It is important to note that different 0-TM sites may share the same Li atoms in an FCC framework and therefore the conversion of different 0-TM to Li thus may result in delithiated structures with varying Li concentrations in Step (2).
 It is more appropriate to use the minimum/maximum voltage as a selection criterion rather than the energy itself for suggesting the next step in the enumeration method.